Consider a population with a mean of 70 and a standard deviation of 6. A random sample of size 36 is drawn. What is the probability that the sample mean is less than 67.5?
1
Expert's answer
2020-05-31T19:16:48-0400

Let X = the population random variable

"\mu" = E(X) = 70 and "\sigma"2 = Var(X) = 62 =36


Let "\bar{x}" be the mean of the random sample of size n taken from this population.


By Central Limit Theorem we know that "\bar{x}" ~ N(μ, σ2/n) asymptotically i.e. for large n.


Therefore, Z = "\frac{(\bar{x}-\mu)}{\sigma/\sqrt{n}}" ~ N(0, 1) asymptotically.


Here, n = 36


Now, the probability that the sample mean is less than 67.5

= P("\bar{x}" < 67.5)

= P("\frac{(\bar{x}-70)}{6/\sqrt{36}}" < "\frac{67.5-70}{6/\sqrt{36}}")

= P(Z < - 2.5)

= "\Phi"(- 2.5)

= 0.0062 [obtained from standard normal distribution table]


Answer: The probability that the sample mean is less than 67.5 is 0.0062.